OpenStudy (anonymous):

A box contains 5 red marbles, 3 blue marbles and 2 yellow marbles. A marble is chosen at random and replaced. This selection process is completed eight times. Find the probability that exactly 4 reds are selected

OpenStudy (anonymous):

Could someone please explain it to me? @satellite73

OpenStudy (anonymous):

what is the probability you get an red marble on any one try?

OpenStudy (mathmate):

Every time the marble is chosen, the probability is the same because marbles are replaced. Therefore use of the binomial distribution is indicated. Have you done binomial distributions?

OpenStudy (anonymous):

1/10

OpenStudy (anonymous):

oh hell no

OpenStudy (anonymous):

sorry 5/10 = 1/2

OpenStudy (anonymous):

would n= 5 and k=4?

OpenStudy (anonymous):

and p=1/2

OpenStudy (anonymous):

ok better so half the time you get it and half the time you don't get it compute \[\binom{8}{4}\left(\frac{1}{2}\right )^{8}\]

OpenStudy (anonymous):

\(n=8, k=4, p=1-p=\frac{1}{2}\)

OpenStudy (anonymous):

so there is no need for bernoulli distribution. what do they exactly mean by this sentence "A marlbe is chosen at random and replaced"?

OpenStudy (anonymous):

these are bernoulli trials and you do use the binomial distribution

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

but since both \(p\) and \(1-p\) are \(\frac{1}{2}\) you don't need to write something silly like \[\binom{8}{4}(\frac{1}{2})^4\times (\frac{1}{2})^4\]

OpenStudy (anonymous):

so when they say "this selection process is completed 8 times" it mean we pick from the box 8 marbles at a time?

OpenStudy (anonymous):

im starting to get it now

OpenStudy (mathmate):

Each pick of the marble is a Bernoulli trial. But when you pick 8 times and want to find 4 success (red), you need the Binomial distribution @satellite73 gave you.

OpenStudy (anonymous):

what if it asked me about the probability of 1 red marble and 2 blue marbles?

OpenStudy (mathmate):

You pick one marble at a time, out of 8 marbles all the time. That is why the probability remains 1/2 for all 8 trials (because of replacement).

OpenStudy (anonymous):

of obtaining*

OpenStudy (anonymous):

ill try do it on my on

OpenStudy (anonymous):

own*

OpenStudy (anonymous):

and correct me later if i am mistaken

OpenStudy (mathmate):

That will be a little more complicated, because binomial distribution does not apply any more when your successful event include more than one colour.

OpenStudy (anonymous):

ok could you please show me how it is done

OpenStudy (mathmate):

Binomial distribution applies only to Bernoulli trials.

OpenStudy (mathmate):

@satellite73 's expression breaks down as follows: \( C^n_r\ p^r\ q^{n-r} \) and applies to n Bernoulli trials with r successes (and n-r failures) with a probability of success equal to p (and probability of failure q=1-p).

OpenStudy (anonymous):

so how is it done if they were to ask me to find the probability of mixed marlbes?

OpenStudy (mathmate):

In the given problem, n=8 (trials) r=4 (successes) n-r=4 (failures) p= 1/2 (probability of success) q= 1-1/2= 1/2 (probability of failure) and \( C^n_r =\frac{n!}{r!(n-r)!} = \frac{8!}{4!4!}\) So the final probability comes down to: \( P(4\ reds) = \frac{8!}{4!4!}(\frac{1}{2})^4(\frac{1}{2})^4 \)

OpenStudy (anonymous):

thanks guys for your help and time. i really appreciate it

OpenStudy (mathmate):

You're welcome! Hope you get the hang of the binomial distribution.

OpenStudy (anonymous):

but what if i was asked to find the probability of mixed marlbes? for instance 1 red and 2 blue?

OpenStudy (anonymous):

how can it be solved

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

would it look like this (10 5) (10 3)/ (10 3)????

OpenStudy (anonymous):

sorry (10 1) (10 2)/ (10 3)????

OpenStudy (anonymous):

im stuffing up real bad :(

OpenStudy (mathmate):

I do not know of an explicit formula for that, see if @satellite73 has a clue. This problem will have to resort to counting, or probability trees. Have you done any of these?

OpenStudy (anonymous):

i guess it is (5 1) (3 2) /(10 3)

OpenStudy (anonymous):

yes with @satellite73

OpenStudy (anonymous):

it should be 0.125

OpenStudy (mathmate):

In this case, we can still use the binomial distribution to pick 5 yellows, namely \( C^8_5\ (\frac{1}{5})^5 \ (\frac{4}{5})^3 \) There are 8 possible arrangements for the remaining marbles, red and blue, namely RRR RRB RBR RBB BRR BRB BBR BBB out of which 3 are 1 red and 2 blues. So multiply the above by \( \frac{3}{8} \) to get \( \frac{3}{8}\ C^8_5\ (\frac{1}{5})^5 \ (\frac{4}{5})^3 \)

OpenStudy (anonymous):

one red and two blue out of how many picked? three?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

out 0f 10 marlbes in total

OpenStudy (anonymous):

you had it above

OpenStudy (anonymous):

that is how you taught me to do it

OpenStudy (anonymous):

is it right?

OpenStudy (anonymous):

\[\frac{\binom{5}{1}\binom{3}{2}}{\binom{10}{3}}\]

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

thanks little by little im getting good at this :)

OpenStudy (anonymous):

yeah practice is good

OpenStudy (mathmate):

@satellite73 just to make sure, the solution is for picking 3 marbles, right?

OpenStudy (anonymous):

yes

OpenStudy (mathmate):

thank you!

OpenStudy (anonymous):

yw clear right? it is instant to write the solution, just takes a minute to compute

OpenStudy (anonymous):

i got an ice chest with 7 miller, 8 bud, 10 strohs if i pick 4 at random, the probability i get 1 miller 2 buds and a stohs is \[\frac{\binom{7}{1}\binom{8}{2}\binom{10}{1}}{\binom{15}{4}}\]

OpenStudy (mathmate):

Do you mean C(25,4) at the bottom?